• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R.A. Lindgren

    Log entry time 16:01:35 on February17,2007

    Entry number 193872

    keyword=Shift Summary

    Riochard Lindgren
    Huan Yao
    Run 6522/26506 In progress BC target 51 deg ended 8.11 AM 1M events
    Run 6523/26507 Start 8:13 AM BC target 51 deg ended 8:41 AM 1M events
    Increase beam current to 1 ua
    Run 6524/26508 Start 8:50 AM Li target 51 deg ended 9:56 AM .5M events
    Increase beam current to 4 ua
    The following 6525/26509 and 6526/26510 end of run halog entries for both HRS are incorrectly labeled for the beam current
    Run 6525/26509 Start 9:57 AM Li target 51 deg ended 11:00 AM 1M events
    Take another 0.5 M events for Li
    Run 6526/26510 Start 11:03AM Li target 51 deg ended 1:35 AM .5M events
    Various replays short and full are in progress by Huan
    Change LHRS angle to 55.5 degrees
    Run 6527/26511 Start 11:50 AM Li target 55.5 deg ended
    Run 6528 Start 13:07 Li Target 55.5 deg ended 14:08 .78M events
    Run 26512 Start 13:07 Li target 55.5 deg end of run failed TS0 supervisor hung up.
    Restart Coda
    Run 26513 Junk
    Run 6530/26514 Prescaler incorrect Junk
    Run 6531/26515 Start 14:41 BC target 55.5 deg ended 15:43 1M events
    Run 6532/26516 Start 15:53 BC target 55.5 deg ended
    Shift Ends